consider the following economy, where prices are fixed: c=100+3/4*( disposable income, t= 1/3*y , i=100, g=300, x=100, m=1/6*y. what is the equation for the aggregate expenditure curve?


1
Expert's answer
2021-08-16T08:57:52-0400

Solution:

The equation for aggregate expenditure curve is: AE = C + I + G + NX

Where: C = Consumption

             I = Investment

            G = Government spending

          NX = Net Exports (Exports – Imports)

T = 0.3Y


C = 100 + 0.75YD = 100 + 0.75 (Y – T) = 100 + 0.75 (Y – 0.3Y)

I = 100

G = 300

NX = x – m = 100 – 0.16Y

AE = 100 + 0.75(Y – 0.3Y) + 100 + 300 + 100 – 0.16Y

AE = 100 + 100 + 300 + 100 + 0.75Y – 0.225Y – 0.16Y

AE = 600 + 0.365Y


The equation for aggregate expenditure curve is: AE = 600 + 0.365Y